# Typical Quantum States of the Universe are Observationally Indistinguishable

### v1

@@
Abstract: This paper is about the epistemology of quantum theory. We establish a new result about a limitation to knowledge of its central object—the quantum state of the universe. We show that, if the universal quantum state can be assumed to be a typical unit vector from a high-dimensional subspace of Hilbert space (such as the subspace defined by a low-entropy macro-state as prescribed by the Past Hypothesis), then no observation can determine (or even just narrow down significantly) which vector it is. Typical state vectors, in other words, are observationally indistinguishable from each other. Our argument is based on a typicality theorem from quantum statistical mechanics. We also discuss how theoretical considerations that go beyond the empirical evidence might bear on this fact and on our knowledge of the universal quantum state.
Key words: limitation to knowledge; empirical equivalence; past hypothesis; quantum statistical mechanics.
@@ 1 Introduction
This paper is about the epistemology of quantum theory. We establish a new result about a limitation to knowledge of the quantum state of the universe. We show that typical universal quantum states are observationally indistinguishable, in the sense that no observation can distinguish typical universal quantum states in a high-dimensional subspace of the Hilbert space, such as the subspace characterized by a low-entropy macro-state as prescribed by the Past Hypothesis (see §2.2 and (Albert 2000, Goldstein et al. 2020)). We call this fact observation typicality.
@@
Observation typicality follows from a result in quantum statistical mechanics called distribution typicality (Reimann 2007, Teufel et al. 2023): for any observable, most quantum states in a high-dimensional Hilbert subspace lead to very nearly the same probability distribution for that observable. We show that this implies observation typicality.
Observation typicality is a surprisingly strong result. There are other known examples of epistemic limitations in physics. Dürr et al. (2004) already proved that it is impossible to measure a wave function precisely. Observation typicality is a further step in this direction, and it delivers a stronger conclusion—even an imprecise measurement is impossible. No observation, as long as it does not disconfirm the physical theory, will give us any substantive information about the quantum state. In this sense, limitation to knowledge is more pervasive and persistent in a quantum universe than has been recognized.
A statement about most quantum states in the full Hilbert space $\mathscr{H}$ of the universe has limited applicability because it may fail for the quantum states of interest. For example, most quantum states in $\mathscr{H}$ are in thermal equilibrium, but the actual quantum state of our universe is not. To use typicality theorems in our situation, we apply them to a subspace $\mathscr{H}_{0}$ of $\mathscr{H}$ such as the subspace comprising the quantum states $\Psi$ compatible with a certain macro-state of the universe that may be far from thermal equilibrium. We assume that $\mathscr{H}_{0}$ has finite but huge dimension $d_{0}$ (while $\mathscr{H}$ may have finite or infinite dimension). Observation typicality then holds for most quantum states $\Psi$ in $\mathscr{H}_{0}$. We can use the result while accepting certain inductive hypotheses about the physical laws of the universe, including those pertaining to initial conditions, according to which our universe is a typical member of the possible universes compatible with such laws.
@@
We assume here that there is a wave function $\Psi_{t}$ of the universe and that at the initial time $t_{0}$ of the universe (say, at the big bang), $\Psi_{t_{0}}$ had to lie in a particular subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe (say, corresponding to a low-entropy macro-state, as prescribed by the Past Hypothesis). Observation typicality asserts that our empirical data at any time $t$ will reveal very little about $\Psi_{t_{0}}$ (beyond its membership in $\mathscr{H}_{0}$, which we take as known) because most $\Psi_{t_{0}}$ from $\mathscr{H}_{0}$ lead to very nearly the same probability distribution for these empirical data. In other words, typical initial quantum states of the universe are observationally indistinguishable from each other. If the time evolution is unitary, then the same arguments yield that also $\Psi_{t}$ cannot be distinguished empirically from typical alternative vectors in the appropriate subspace, i.e., in $U_{t}\mathscr{H}_{0}$ with $U_{t}=\exp(-iH(t-t_{0}))$ and $H$ the Hamiltonian of the universe.
@@
The analogous statement fails in classical mechanics. Observation typicality holds in quantum mechanics because, roughly speaking, quantum states can be in superpositions, and the superposition weights of any particular observation given by typical quantum states are nearly the same. The latter is related to the mathematical fact, known as the concentration of measure phenomenon (the situation where random quantities become nearly deterministic), which applies in any high-dimensional Hilbert space.
@@
We present two main results. First (§2 and §3), that a typical unit vector $\Psi_{t_{0}}\in\mathscr{H}_{0}$ cannot be reliably distinguished from the density matrix $\rho_{0}$ associated with a uniform distribution in $\mathscr{H}_{0}$, i.e., $\rho_{0}=P_{0}/d_{0}$ with $P_{0}$ the projection operator to $\mathscr{H}_{0}$ and $d_{0}=\dim\mathscr{H}_{0}$; as a consequence, typical unit vectors cannot be reliably distinguished from each other. The second result (§3.4) is even stronger and asserts that even unreliably, they cannot be distinguished; that is, our empirical data do not even yield (any significant amount of) partial or probabilistic information about $\Psi_{t_{0}}$. This result can be expressed in terms of Bayesian credences as follows: if we start from the uniform probability distribution $u_{0}$ over the unit sphere $\mathbb{S}(\mathscr{H}_{0})$ (containing all normalized wave functions) in $\mathscr{H}_{0}$ as the prior distribution, and if we update our credences in a Bayesian way on the basis of our empirical observation (that is not too improbable), then the updated distribution will still be very nearly uniform over $\mathbb{S}(\mathscr{H}_{0})$.
@@
Some empirical observations might, of course, change our mind about which subspace $\mathscr{H}_{0}$ is the correct one. Since we regard here the specification of $\mathscr{H}_{0}$ as part of the given theory, such an observation would amount to a disconfirmation of the theory. This leads to the interesting question of which kind of observations should be regarded as disconfirming the theory. We will look more into this question elsewhere (Chen and Tumulka 2024), take here for granted that the theory is correct and not getting disconfirmed, and remark only briefly that since the specification of $\mathscr{H}_{0}$ has the status of a law of nature, we may expect it to be simple and natural in order to be convincing; as a consequence, we would not be inclined to adjust the choice of $\mathscr{H}_{0}$ in complicated ways just to adapt it to some random-looking features of our world (such as, e.g., the shape of Ireland).
@@
Our goal is to further the discussion about the epistemology of physical theories and highlight the limitations of (direct) empirical knowledge in quantum universes. Insofar as we have detailed knowledge about the universal quantum state, it is even more theoretical than has been recognized. Our result should be of interest to anyone interested in the epistemological implications of physics. We highlight several distinctive features of the present approach:
  • Observation typicality applies to all interpretations of quantum mechanics. It is also compatible with generous inductive assumptions about what we can infer based on available evidence. For example, it is compatible with our knowledge of the dynamical laws of quantum mechanics, the Past Hypothesis, and the aggregation of individual observations into a collection.
  • Our argument relies on the concentration of measure phenomenon instead of special constructive techniques, such as the cut-and-paste method used in Manchak (2009)’s argument for observational indistinguishability in general relativity. The standard “physical unreasonableness” objections against arguments based on special constructions do not apply to our argument, which establishes observational indistinguishability for generic pairs of models.
  • Our claims are based on quantitative bounds (recently proven mathematically) on how large the deviations of certain probabilities can at most become for which fraction of all wave functions in a high-dimensional Hilbert subspace.
@@
In §2, we formulate distribution typicality and explain how it applies to the special case of the Past Hypothesis. In §3, we clarify our criterion for observational indistinguishability, derive observation typicality from distribution typicality, apply the fact to the universal quantum state, and provide a Bayesian analysis. In §4, we compare observation typicality to known results about epistemic limitations in physics. In §5, we discuss potential philosophical implications. For more detail, see the longer version, Chen and Tumulka (2024).
@@ 2 Distribution Typicality
In the 21st century, many exciting new results have been proven in quantum statistical mechanics and quantum thermodynamics, leading to improved understanding in the thermodynamic behavior of a closed quantum system in a pure state (Gemmer et al. 2009, Tasaki 2016, Mori et al. 2018). Such results often assume an “individualist” attitude, according to which an individual quantum system in a pure state, as opposed to an ensemble, is studied with respect to its thermodynamic properties. These results often concern typical properties of those states. [^3] There is a growing philosophical literature about typicality and its role in physical theories. See, for example, Goldstein (2012), Reichert (2018), Wilhelm (2022), Lazarovici (2023), Hubert (2024), (Frigg and Werndl 2024, sec. 4.5), and the references therein.
We shall focus on distribution typicality in this paper. In §3, we show that distribution typicality implies observation typicality. In this section, we formulate distribution typicality and specialize it to the Past Hypothesis.
@@ 2.1 The Key Theorem
Before stating the key theorem, we highlight its implications for orthodox quantum mechanics (OQM), Bohmian mechanics (BM), and the Ghirardi-Rimini-Weber spontaneous collapse theory (GRW) (Tumulka 2022b, Dürr and Teufel 2009, Ghirardi and Bassi 2020). Distribution typicality entails that, in OQM, if an observer at time $t$ conducts an experiment, the probability distribution of the outcome is nearly independent of $\Psi_{t_{0}}$; in BM, at any time $t$, the probability distribution of the Bohmian configurations $Q_{t}$ of the universe is nearly independent of $\Psi_{t_{0}}$; in GRW with a flash ontology, for any time $t>t_{0}$, the probability distribution of the pattern of flashes up to $t$ is nearly independent of $\Psi_{t_{0}}$. We will elucidate in §3 why this near-independence entails the impossibility for observers according to OQM, BM, or GRW to distinguish among different $\Psi_{t_{0}}$.
Terminology. Recall that $\mathbb{S}(\mathscr{H}_{0})$ is the unit sphere in the subspace $\mathscr{H}_{0}$, $d_{0}=\dim\mathscr{H}_{0}$, and $u_{0}$ the uniform probability distribution over $\mathbb{S}(\mathscr{H}_{0})$; that is, for any subset $S$ of $\mathbb{S}(\mathscr{H}_{0})$, $u_{0}(S)$ is the surface area of $S$, normalized by dividing through the surface area of $\mathbb{S}(\mathscr{H}_{0})$. We say that a statement $s(\Psi)$ is true for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ if and only if the set $S$ of $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ satisfying $s(\Psi)$ has $u_{0}(S)\geq 1-\varepsilon$ (equivalently, if $S$ has at least the fraction $1-\varepsilon$ of the surface area, or if a purely random point on the sphere has probability $\geq 1-\varepsilon$ to lie in $S$). Averages on $\mathbb{S}(\mathscr{H}_{0})$ will also be taken with respect to $u_{0}$. [^4] The results in this paper can be extended to the Gaussian adjusted projected (GAP) measures introduced by Goldstein et al. (2006b), as is done by Teufel et al. (2024). But here, we consider only the uniform measure $u_{0}$. We also need the concept of POVM.
POVM. POVM stands for positive-operator-valued measure, which is a generalization of the concept of an observable as given by a self-adjoint operator. Technically, a POVM on a Hilbert space $\mathscr{H}$ is a family of positive operators $E_{z}$ on $\mathscr{H}$ that add up to the identity operator. POVMs have an important representational role in quantum mechanics, given by the Main Theorem about POVMs: for every quantum physical experiment $\mathscr{E}$ on a quantum system $\mathscr{S}$ whose possible outcomes lie in a space $\mathscr{Z}$, there exists a POVM E on $\mathscr{Z}$ such that, whenever $\mathscr{S}$ has wave function $\Psi$ at the beginning of $\mathscr{E}$, the random outcome $Z$ has probability distribution given by
@@
See Dürr et al. (2004) and Tumulka (2022b) for proofs.
@@
Let $\mathscr{H}_{0}$ be an arbitrary subspace with finite dimension $d_{0}$ in $\mathscr{H}$, let the density matrix $\rho_{0}=P_{0}/d_{0}$ be the normalized projection to $\mathscr{H}_{0}$, and let the operator $E_{z}$ be an element of an arbitrary POVM on $\mathscr{H}$. Then for every $\varepsilon>0$, for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$,
@@
$\Bigl{|}\langle\Psi|\,E_{z}\,|\Psi\rangle-\operatorname{tr}(\rho_{0}\,E_{z})\Bigr{|}\leq\frac{1}{\sqrt{\varepsilon d_{0}}}\,.$ (2)
@@
Theorem 1 is a particular case of Theorem 3 proven by Teufel et al. (2023). It can also be derived from a theorem proven by Reimann (2007).
It can be shown that $\operatorname{tr}(\rho_{0}E_{z})$ is exactly the average of $\langle\Psi|E_{z}|\Psi\rangle$ over $\mathbb{S}(\mathscr{H}_{0})$ (using the uniform distribution $u_{0}$); thus (2) expresses that $\langle\Psi|E_{z}|\Psi\rangle$ is close to its average value, provided that the right-hand side is small, which will be the case whenever
$d_{0}\gg\frac{1}{\varepsilon}\,.$ (3)
@@
Inequality (2) is an example of the “concentration of measure phenomenon,” which can be expressed as saying that some random quantity is nearly deterministic (i.e., very probably close to a deterministic value). Here, we have that for any observation represented by an element in a POVM, typical quantum states in a high-dimensional Hilbert space have very nearly the same Born probability distribution for that observation.
Let us clarify and elaborate on this result.
Example. A macroscopic system with $N=10^{20}$ particles is in macro-state $\mu$. The dimension of the Hilbert space $\mathscr{H}_{0}$ of the system scales exponentially with $N$. Let us suppose $d_{0}=10^{10^{20}}$. Consider a very small error $\varepsilon=10^{-200}$. Inequality (2) tells us that, for any quantum experiment with observational outcome $z$, for $(1-\varepsilon)$-most pure states in $\mathscr{H}_{0}$, the Born probability of outcome $z$ given by the pure state differs from the ensemble average by at most $10^{100-0.5\times 10^{20}}\approx 3^{-10^{20}}$, which is close to zero.
@@
Minimal Assumptions. Theorem 1 holds in great generality. First, it applies to every POVM. By the Main Theorem about POVMs, the theorem covers every observational outcome arising from an arbitrary quantum physical experiment. Second, it does not make assumptions about the interaction between the subsystem and the environment. Third, it does not invoke chaos, ergodicity, or mixing. The typicality in distribution typicality comes from the typicality of quantum states in $\mathbb{S}(\mathscr{H}_{0})$, and the large number comes from $d_{0}$, the large dimension of $\mathscr{H}_{0}$. Finally, distribution typicality applies to thermal equilibrium and non-equilibrium. For concreteness, in §2.2 we provide a physical interpretation for a universe in thermal non-equilibrium, taking $\mathscr{H}_{0}$ to be the low-entropy initial macro-state of the universe prescribed by the Past Hypothesis.
@@
Difference from the Classical Case. We do not have distribution typicality in classical mechanics. The prediction of any observation from a particular micro-state in the classical phase space has a trivial probability of 00 or $1$, while the expectation values for many observables with respect to a uniform probability distribution (the ensemble averages) are often non-trivial. It is false that for every experimental outcome, typical micro-states will assign very nearly the same probability distribution. For example, we can design a coin-flip experiment where half the microstates assign probability 00 and the other half assign probability $1$ to the “Tails” outcome. The relevant way in which quantum mechanics differs from classical mechanics is that if a wave function $\Psi=\Psi_{t_{0}}$ gets chosen randomly, then the outcome $Z$ of an observation is “doubly random”: Given $\Psi$, it is random with the Born distribution $\langle\Psi|E_{z}|\Psi\rangle$, and in addition $\Psi$ is itself random. The crucial point is that in our situation, different $\Psi$s have very similar superposition weights, with the consequence that they have nearly the same $\langle\Psi|E_{z}|\Psi\rangle$.

### v2

@@
Abstract: We establish three new impossibility results regarding our knowledge of the quantum state of the universe—a central object in quantum theory. We show that, if the universal quantum state is a typical unit vector from a high-dimensional subspace $\mathscr{H}_{0}$ of Hilbert space $\mathscr{H}$ (such as the one defined by a low-entropy macro-state as prescribed by the Past Hypothesis), then no observation can determine or just significantly narrow down which vector it is. In other words, the overwhelming majority of possible state vectors are observationally indistinguishable from each other (and from the density matrix of $\mathscr{H}_{0}$). Moreover, we show that for any observation that isn’t too unlikely and most pairs of unit vectors from $\mathscr{H}_{0}$, the observation will not significantly favor one vector over the other. We further show that the uniform distribution over the unit sphere in $\mathscr{H}_{0}$, after Bayesian updating in the light of any observation that isn’t too unlikely, is still extremely close to uniform. Our arguments rely on a typicality theorem from quantum statistical mechanics. We also discuss how theoretical considerations beyond empirical evidence might inform our understanding of this fact and our knowledge of the universal quantum state.
Key words: limitation to knowledge; empirical equivalence; past hypothesis; quantum statistical mechanics; quantum measurement; measure concentration.
@@ 1 Introduction
We establish three new impossibility results regarding our knowledge of the quantum state of the universe—a central object in quantum theory. We show that typical universal quantum states are observationally indistinguishable, in the sense that no observation can distinguish typical universal quantum states in a high-dimensional subspace of the Hilbert space, such as the one prescribed by the Past Hypothesis (see §2.2 and (Albert 2000, Goldstein et al. 2020)), unless we somehow already knew the quantum state of the universe. We call this fact observation typicality.
@@
Observation typicality follows from a result in quantum statistical mechanics called distribution typicality (Reimann 2007, Teufel et al. 2023): for any observable, most quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$ lead to very nearly identical probability distributions for that observable. Roughly speaking, that is, most wave functions from a high-dimensional $\mathscr{H}_{0}$ look the same.
@@
We assume here that there is a wave function $\Psi_{t}$ of the universe and that at the initial time $t_{0}$ of the universe (say, at the big bang), $\Psi_{t_{0}}$ had to lie in a particular subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe (say, corresponding to a low-entropy macro-state, §2.2). Observation typicality then asserts that our empirical data at any time $t$ will reveal very little about $\Psi_{t_{0}}$ (beyond its membership in $\mathscr{H}_{0}$, which we take as known), even if we assume that any observable (including any POVM) can be measured.
Observation typicality is a surprisingly strong result, yet it has gone unnoticed so far. There are other known examples of epistemic limitations in physics. Dürr et al. (2004) already proved that it is impossible to measure a wave function precisely. Observation typicality takes this limitation a step further, delivering an even stronger conclusion: typically, even an imprecise measurement is impossible. No observation—provided it does not disconfirm the physical theory—can yield substantive information about the quantum state. In this sense, the limitation to knowledge is more pervasive and persistent in a quantum universe than previously recognized.
We note that a statement about most quantum states in the full Hilbert space $\mathscr{H}$ of the universe would have limited applicability because it may fail for the quantum states of interest. For example, most quantum states in $\mathscr{H}$ are in thermal equilibrium, but the actual quantum state of our universe is not. To use typicality theorems in our situation, we apply them to a subspace $\mathscr{H}_{0}$ of $\mathscr{H}$ such as the subspace comprising the quantum states $\Psi$ compatible with a certain macro-state of the universe that may be far from thermal equilibrium. We assume that $\mathscr{H}_{0}$ has finite but huge dimension $d_{0}$ (while $\mathscr{H}$ may have finite or infinite dimension). [^3] If our universe initially had a finite volume and occupied a finite energy interval $[E_{1},E_{2}]$, it is confined to a Hilbert space $\mathscr{H}_{[E_{1},E_{2}]}$ of finite (albeit realistically very large) dimension. See §2.2. We can use the result while accepting certain physical laws, including those pertaining to initial conditions of the universe, according to which our universe is a typical member of the possible universes compatible with such laws.
It is well known that if we pick two vectors $\Psi_{A},\Psi_{B}$ purely randomly from the unit sphere of a high-dimensional $\mathscr{H}_{0}$, then with overwhelming probability they will be nearly orthogonal, and thus there exists a quantum observable $P$ that distinguishes them reliably. However, this fact does not affect the question we are studying here because the choice of $P$ depends on (at least one of) $\Psi_{A},\Psi_{B}$, while we do not assume we already know the wave function of the universe. The situation that we do not allow the observable to depend on the quantum state is reflected in formal statements by the order of quantifiers such as some, most, and every. To clarify the order of quantifiers in our result: By observation typicality, we mean that for every observation, most quantum states are not distinguishable by that observation. The alternative order—where we first quantify over most states and then over every observation—would imply an even stronger claim, which we call super-strong observation typicality. The latter claim is generally false, though it holds for local observations because of the more well-known result of canonical typicality, as we discuss in Chen and Tumulka (2025).
@@
In classical mechanics, the statement analogous to observation typicality fails. It holds in quantum mechanics because, roughly speaking, quantum states can be in superpositions, and the superposition weights of any particular observation given by typical quantum states are nearly identical. The latter is a version of the mathematical fact known as the concentration-of-measure phenomenon (the situation where random quantities become nearly deterministic) and occurs in high-dimensional Hilbert spaces. We note further that the same arguments used to establish observation typicality yield that, if the time evolution is unitary, also $\Psi_{t}$ (instead of $\Psi_{t_{0}}$) cannot be distinguished empirically from typical alternative vectors in the appropriate subspace, i.e., in $U_{t}\mathscr{H}_{0}$ with $U_{t}=\exp(-iH(t-t_{0}))$ and $H$ the Hamiltonian of the universe.
@@
Our conclusions are based on three mathematical results. First (§3.2.1), that a typical unit vector $\Psi_{t_{0}}\in\mathscr{H}_{0}$ cannot be reliably distinguished from the density matrix $\rho_{0}$ associated with a uniform distribution in $\mathscr{H}_{0}$, i.e., $\rho_{0}=P_{0}/d_{0}$ with $P_{0}$ the projection operator to $\mathscr{H}_{0}$ and $d_{0}=\dim\mathscr{H}_{0}$. Consequently, typical unit vectors cannot be reliably distinguished from each other. By contrast, the second result (§3.2.2) is in a sense stronger but does require that the observation in question be not too improbable. It asserts that even unreliably, the states cannot be distinguished, which conveys that such empirical data do not even yield (any significant amount of) partial or probabilistic information about $\Psi_{t_{0}}$. The third result (§3.2.3) is expressed in Bayesian terms: if we start from the uniform prior $u_{0}$ over the unit sphere $\mathbb{S}(\mathscr{H}_{0})$ (containing all normalized wave functions) in $\mathscr{H}_{0}$ and update in a Bayesian way upon an observation that is not too improbable, then the posterior distribution will still be very nearly uniform over $\mathbb{S}(\mathscr{H}_{0})$.
@@
Some empirical observations might, of course, change our mind about which subspace $\mathscr{H}_{0}$ is the correct one. Since we regard here the specification of $\mathscr{H}_{0}$ (in the Past Hypothesis) as part of the given theory, such an observation would amount to a disconfirmation of the theory. This leads to the interesting question of which kind of observations should be regarded as disconfirming the theory. We will look more into this question elsewhere (Chen and Tumulka 2025), take here for granted that the theory is correct and not getting disconfirmed, and remark only briefly that if the specification of $\mathscr{H}_{0}$ has the status of a law of nature, we may expect it to be simple and natural in order to be convincing (Loewer 2024, Chen 2024c); as a consequence, we would not be inclined to adjust the choice of $\mathscr{H}_{0}$ in complicated ways just to adapt it to some random-looking features of our world (such as, e.g., the shape of Ireland).
@@
Our results show that insofar as we have detailed knowledge about the universal quantum state, it is even more theoretical than has been recognized. They should be of interest to anyone interested in the epistemological implications of physics.
@@
This paper is organized as follows. In §2, we formulate distribution typicality and explain how it applies to the special case of the Past Hypothesis. In §3, we clarify our criterion for observational indistinguishability, derive observation typicality from distribution typicality, and provide a Bayesian analysis. In §4, we compare observation typicality to known results about epistemic limitations in physics. In §5, we discuss potential philosophical implications. For more detail, including the application of observation typicality to subsystem quantum states of the universe, see Chen and Tumulka (2025). In §6, we conclude.
@@ 2 Distribution Typicality
In the 21st century, many exciting new results have been proven in quantum statistical mechanics and quantum thermodynamics, leading to improved understanding of the thermodynamic behavior of a closed macroscopic quantum system in a pure state (Gemmer et al. 2009, Tasaki 2016, Mori et al. 2018). Such results often assume an “individualist” attitude, according to which an individual quantum system in a pure state, as opposed to either a density matrix or an ensemble of systems, is studied with respect to its thermodynamic properties. These results often concern typical properties of those states. [^4] The concept of typicality and its role in physical theories have received growing attention in the philosophical literature. For some recent examples, see Goldstein (2012), Lazarovici and Reichert (2015), Allori (2020), Wilhelm (2022), Lazarovici (2023), Hubert (2024), (Frigg and Werndl 2024, sec. 4.5), and the references therein.
We shall focus on distribution typicality in this paper. In §3, we show that distribution typicality implies observation typicality for orthodox quantum mechanics (OQM), Bohmian mechanics (BM), the Ghirardi-Rimini-Weber spontaneous collapse theory (GRW), and Everettian quantum mechanics (EQM) (Tumulka 2022b, Dürr and Teufel 2009, Ghirardi and Bassi 2020, Barrett 2023). In this section, we formulate distribution typicality and specialize it to the Past Hypothesis.
@@ 2.1 The Key Theorem
Terminology. Recall that $\mathbb{S}(\mathscr{H}_{0})$ is the unit sphere in the subspace $\mathscr{H}_{0}$, $d_{0}=\dim\mathscr{H}_{0}$, and $u_{0}$ the uniform probability distribution over $\mathbb{S}(\mathscr{H}_{0})$; that is, for any subset $S$ of $\mathbb{S}(\mathscr{H}_{0})$, $u_{0}(S)$ is the surface area of $S$, normalized by dividing through the surface area of $\mathbb{S}(\mathscr{H}_{0})$. We say that some statement about the wave function $\Psi$, $s(\Psi)$, is true for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ if and only if the set $S$ of $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ satisfying $s(\Psi)$ has $u_{0}(S)\geq 1-\varepsilon$ (equivalently, if $S$ has at least the fraction $1-\varepsilon$ of the surface area, or if a purely random point on the sphere has probability $\geq 1-\varepsilon$ to lie in $S$). Averages on $\mathbb{S}(\mathscr{H}_{0})$ will also be taken with respect to $u_{0}$. [^5] The results in this paper can be extended to the Gaussian adjusted projected (GAP) measures introduced by Goldstein et al. (2006b), as is done by Teufel et al. (2025). But here, we consider only the uniform measure $u_{0}$. We also need the concept of POVM.
POVM. A POVM (positive-operator-valued measure) is a generalization of the notion of an observable represented by a self-adjoint operator. In the simple case with discrete outcomes, a POVM on a Hilbert space $\mathscr{H}$ is a family of positive operators $E_{z}$ on $\mathscr{H}$ that add up to the identity operator. More generally, a POVM is a map $E$ from the measurable subsets of an outcome space $\mathscr{Z}$ to the set of positive operators on $\mathscr{H}$ satisfying normalization, $E(\mathscr{Z})=I$, and countable additivity. POVMs have an important representational role in quantum mechanics, given by the Main Theorem about POVMs: for every quantum physical experiment $\mathscr{E}$ on a quantum system $\mathscr{S}$ whose possible outcomes lie in a space $\mathscr{Z}$, there exists a POVM E on $\mathscr{Z}$ such that, whenever $\mathscr{S}$ has wave function $\Psi$ at the beginning of $\mathscr{E}$, the random outcome $Z$ has probability distribution given by
@@
where $E_{z}$, the image of $\{z\}$ under $E$, is an element of the POVM $E$. See Dürr et al. (2004) and Tumulka (2022b) for proofs.
@@
Let $\mathscr{H}_{0}$ be an arbitrary subspace with finite dimension $d_{0}$ in $\mathscr{H}$, let the density matrix $\rho_{0}=P_{0}/d_{0}$ be the normalized projection to $\mathscr{H}_{0}$, and let $B$ be any self-adjoint operator on $\mathscr{H}$. Then for every $\varepsilon>0$, for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$,
@@
$\Bigl{|}\langle\Psi|\,B\,|\Psi\rangle-\operatorname{tr}(\rho_{0}\,B)\Bigr{|}\leq\frac{\|B\|}{\sqrt{\varepsilon d_{0}}}\,,$ (2)
where the operator norm $\|B\|$ means the largest absolute eigenvalue of $B$. Moreover, for every POVM $\{E_{1},\ldots,E_{L}\}$ and every $\varepsilon>0$, for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$, the distribution $\langle\Psi|E_{z}|\Psi\rangle$ is close to the distribution $\operatorname{tr}(\rho_{0}\,E_{z})$ in the sense that for each $z$,
$\Bigl{|}\langle\Psi|\,E_{z}\,|\Psi\rangle-\operatorname{tr}(\rho_{0}\,E_{z})\Bigr{|}\leq\sqrt{\frac{L}{\varepsilon d_{0}}}\,.$ (3)
Note that the right-hand side of (3) is small whenever
$d_{0}\gg\frac{L}{\varepsilon}\,.$ (4)
@@
Theorem 1 is a particular case of Theorem 3 proven by Teufel et al. (2023). It can also be derived from a theorem proven by Reimann (2007). The statement involving Eq. (3) follows from the one involving (2). [^6] Here is how: Replace $B$ by $E_{z}$ and $\varepsilon$ by $\varepsilon/L$ in the statement involving (2) and use $\|E_{z}\|\leq 1$. Thus, for any $z$, the set of $\Psi$’s satisfying (3) has measure at least $1-\varepsilon/L$, so the intersection of these sets has measure at least $1-\varepsilon$.
@@
It can be shown that $\operatorname{tr}(\rho_{0}E_{z})$ is exactly the average of $\langle\Psi|E_{z}|\Psi\rangle$ over $\mathbb{S}(\mathscr{H}_{0})$ (using the uniform distribution $u_{0}$); thus, (3) expresses that $\langle\Psi|E_{z}|\Psi\rangle$ is close to its average value whenever (4) holds. Theorem 1 is an example of the “concentration-of-measure phenomenon.” Here, we have that for any observation represented by a POVM, typical quantum states in a high-dimensional Hilbert space have very nearly identical Born probability for that observation. Let us clarify and elaborate on this theorem.
Example 1.
Which size should we consider for $d_{0}$ and $\varepsilon$ in practice? For a macroscopic quantum system with $N$ particles, the dimension of an energy shell is usually of order $10^{N}$ (or perhaps more like $10^{1000\,N}$). While the dimensions of subspaces corresponding to different macro states differ by huge factors, even the comparatively small subspaces of this kind, corresponding to states far from thermal equilibrium, have dimension of order roughly $10^{N}$. For the universe, a widespread estimate is $N=10^{80}$, which motivates us to consider,
[…]
Let us turn to $\varepsilon$. Borel (1962) famously argued that an event with probability,
[…]
or smaller can be assumed to never occur in the history of the universe. Of course, we can make sure that some event of smaller probability occurs, for example by tossing a fair coin 1000 times; then the actual sequence of heads and tails had probability $2^{-1000}\approx 10^{-300}$ like every sequence; but Borel meant that any event specified in advance with probability $\leq 10^{-200}$ is not going to occur. Correspondingly, for this value of $\varepsilon$ and any given POVM, we can expect $\Psi_{t_{0}}$ to lie in the set of $\Psi$’s satisfying (3). While our results do not require the values (5), (6), we will use them for illustration. So for our purposes, (4) is satisfied, for example for $L=2$ or $L=1000$, with enormous margins. Our results in §3 will be based on (2), as it provides the best bound for a given outcome $z$.
Example 2.
[…]
@@
Minimal Assumptions. Theorem 1 holds in great generality. First, it applies to an arbitrary POVM. (In fact, it even applies to every bounded operator.) By the Main Theorem about POVMs, the theorem covers every possible outcome arising from an arbitrary quantum physical experiment. Second, it does not make assumptions about the interaction between the subsystem and the environment. Third, it does not invoke chaos, ergodicity, or mixing. The typicality in distribution typicality comes from the typicality of quantum states in $\mathbb{S}(\mathscr{H}_{0})$ for a given observation, and the large number comes from $d_{0}$, the large dimension of $\mathscr{H}_{0}$. Finally, distribution typicality applies to thermal equilibrium and non-equilibrium. For concreteness, in §2.2 we provide a physical interpretation for a universe in thermal non-equilibrium, taking $\mathscr{H}_{0}$ to be the low-entropy initial macro-state of the universe prescribed by the Past Hypothesis.
@@
Difference from the Classical Case. We do not have distribution typicality in classical mechanics. The prediction of any observation from a particular micro-state in the classical phase space has a trivial probability of 00 or $1$, while for many observables their probability distribution induced by the uniform probability distribution in a region of phase space (the ensemble statistics) is non-trivial. For example, we can design a coin-flip experiment where half the microstates assign probability 00 and the other half assign probability $1$ to the “Tails” outcome. The relevant way in which quantum mechanics differs from classical mechanics is that if a wave function $\Psi=\Psi_{t_{0}}$ gets chosen randomly, then the outcome $Z$ of an observation is “doubly random”: Given $\Psi$, it is random with the Born distribution $\langle\Psi|E_{z}|\Psi\rangle$, and in addition $\Psi$ is itself random. The crucial point is that in our situation, different $\Psi$s have very similar superposition weights, with the consequence that they have nearly identical $\langle\Psi|E_{z}|\Psi\rangle$.